OpenStudy (anonymous):

Given: x – 2(x + 2 + 2x) = 10 + 2x Prove: x = –2 Mathematical Statement Justification x – 2(x + 2 + 2x) = 10 + 2x Given x – 2(x + 2x + 2) = 10 + 2x Commutative Property of Addition x – 2(3x + 2) = 10 + 2x Combine Like Terms x – 6x – 4 = 10 + 2x Distributive Property – 5x – 4 = 10 + 2x – 7x – 4 = 10 – 7x = 14 x= –2 Division of Property Equality What are the missing justification steps, in order? Distributive Property, Addition Property of Equality, Subtraction Property of Equality Combine Like Terms, Subtraction Property of Equality, Addition Property of Equality

OpenStudy (anonymous):

The answers: 1. Distributive Property, Addition Property of Equality, Subtraction Property of Equality 2. Combine Like Terms, Subtraction Property of Equality, Addition Property of Equality 3. Distributive Property, Combine Like Terms, Subtraction Property of Equality 4. Combine Like Terms, Distributive Property, Addition Property of Equality

OpenStudy (anonymous):

Mind helping me?

OpenStudy (anonymous):

well what do they do with the x's in the first part

OpenStudy (anonymous):

distributive property?

OpenStudy (anonymous):

you can see that x-6x and then in the next it is -5x

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

distributive property is this : a(b+d)=ab+ad

OpenStudy (anonymous):

oh is it combine like properties?

OpenStudy (anonymous):

I mean terms

OpenStudy (anonymous):

combine like terms, yes it is taking the x terms and making them one

OpenStudy (anonymous):

alright I got that so far

OpenStudy (anonymous):

subtraction property next?

OpenStudy (anonymous):

yes because the 2x is positive on the right side and disappears in the next line the subtraction property of equality says that you can subtract something from both sides and the equation will be equal still

OpenStudy (anonymous):

So is the answer 2?

OpenStudy (anonymous):

Don't see how it can be 4

OpenStudy (anonymous):

yes lastly the addition property is saying that if you add the same number or term to each side, the answer will be the same... in this case they add 4 to both sides to get rid of the 4 on the left side

OpenStudy (anonymous):

Mind helping me with one more?

OpenStudy (anonymous):

sure post it

OpenStudy (anonymous):

I can just put it in here right?

OpenStudy (anonymous):

Nathan wrote the following steps to prove the equation: Given: 3(x + 2) = 6x + 3 Prove: x = 1 Step Mathematical Statement Justification 0 3(x + 2) = 6x + 3 Given 1 3x + 6 = 6x + 3 Distributive Property 2 6 = 3x + 3 Commutative Property of Addition 3 3x + 3 = 6 Symmetric Property of Equality 4 3x = 3 Subtraction Property of Equality 5 x =1 Division Property of Equality Which step of justification is incorrect, and what should the justification for that step be to solve the equation? Step 4; Commutative Property of Multiplication Step 1; Associative Property of Addition Step 2; Subtraction Property of Equality Step 3; Reflexive Property of Equality